OpenStudy (anonymous):

Find a formula for the inverse of the function. y = ln(x + 5) Please explain

OpenStudy (anonymous):

this says 1) add 5 2) take the log to undo that you will 1) exponentiate 2) subtract 5 it will look like \[f^{-1}(x)=e^x-5\]

OpenStudy (anonymous):

or if you prefer you can write \[x=\ln(y+5)\] and solve for y, but the steps are the same \[x=\ln(y+5)\] \[e^x=y+5\] \[e^x-5=y\] \[f^{-1}(x)=e^x-5\]